28. August 2015: Friday

LDAP/IdM:

  • LDAP Update:
    • A one item match: A squabble of whether it is two last names or a hyphenated last name. ¬†Merged.
    • Code for one-item-match comparisons updated to strip out spaces and hyphens before comparing – since people like to use them in place of each other.
  • ACCOUNT Update:
    • no issues.
  • VPN Update:
    • Thirteen (13) changes made – all appropriate.

Backups:

  • Save Groups:
    • Failed: windowpane – as usual.
  • Clones:
    • All finished.
  • Problems:
    • A0945SG: UPS site says both packages are on the truck to be delivered today.
      • IBM called at 9:48am to say the parts would be here within the hour. ¬†As of 10:55am – no parts.
    • Qualstar claims all tape drives have stopped talking.
      • Emailed Qualstar with the problem description.
      • Called their after-hours support line at 8:00am, got Bryan out of bed (5am his time) and he solved the problem.
        • Verified the physical hardware was all still ok, just communications was failed.
        • stop and start the “drm” process
  • General:
    • Tape schlepping:
      • 168 to go, 168 to come back.

 

Tags:

Leave a Reply

You must be logged in to post a comment.